OpenStudy (anonymous):

HELP PLEASE!! MEDALS WILL BE GIVEN!! Solve the system of equations: 9a+3b+c=2 a+b+c=4 6a+b=0

OpenStudy (anonymous):

replace \(b\) by \(-6a\) in the first two equations

OpenStudy (anonymous):

\[9a-18a+c=2\\a-6a+c=4\]or \[-9a+c=2\\-5a+c=4\]

OpenStudy (anonymous):

alright i got it from here thank you :)

OpenStudy (anonymous):

yw